OpenStudy (anonymous):

A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

i bet it is \[|x − 0.525| \leq 0.01\] am i right?

OpenStudy (anonymous):

lol i think so

OpenStudy (anonymous):

that means that x has to be within \(.01\) units of \(0.525\) either to the right or left you can solve this either in your head, or by writing \[-0.01\leq x-0.525\leq 0.01\] and adding \(0.0525\) all the way across

OpenStudy (anonymous):

thanks :)